### v1

# A Sparse Graph Formulation for Efficient Spectral Image Segmentation

Abstract
Spectral Clustering is one of the most traditional methods to solve segmentation problems. Based on Normalized Cuts, it aims at partitioning an image using an objective function defined by a graph. Despite their mathematical attractiveness, spectral approaches are traditionally neglected by the scientific community due to their practical issues and underperformance. In this paper, we adopt a sparse graph formulation based on the inclusion of extra nodes to a simple grid graph. While the grid encodes the pixel spatial disposition, the extra nodes account for the pixel color data. Applying the original Normalized Cuts algorithm to this graph leads to a simple and scalable method for spectral image segmentation, with an interpretable solution. Our experiments also demonstrate that our proposed methodology over performs traditional spectral algorithms for segmentation.
Index Terms—  Image Segmentation, Spectral Clustering, Sparse Matrices, Optimization.
1 Introduction
Image Segmentation refers to the problem of grouping image pixels in meaningful regions with the goal of understanding or summarizing the visual content in that image [1]. Because of that, image segmentation has found applications in many domains, such as medical image analysis, autonomous driving, and photo editing to name a few [1]. This intense industrial appeal led to the development of many algorithms, which tackled the problem using variational [2], graphical [3, 4, 5], statistical [6] and, more recently, deep leaning techniques [7].
Spectral clustering [8], pioneered by Normalized Cuts [3], provides an algorithm for image segmentation that stands out for its methodological consistency backed by graph theory and computational linear algebra. It consists of encoding the image data into a graph then using spectral techniques to approximately find the cut corresponding to the final segmentation. While popular in its early years, the initial formulation of Normalized Cuts had that would hinder its broad adoption by the scientific community. The issues found were mostly (1) computational, as it required the construction and decomposition of large matrices [9], and (2) practical, as it performed worse than other algorithms [10].
Many strategies were approached to overcome the issues [9, 11, 12, 13]. In this paper, we approach them by adopting a sparse graph consisting of a grid graph that encodes the pixel spatial locations and of a few extra nodes corresponding to the pixel intensities [2, 4]. We show that the Normalized Cut solution found using this graph encourages pixel groupings that linearly balances spatial coherence (neighbors generally belong to the same segment) and color cohesion (pixels in a segment have colors in the same subset of available colors) [14]. Furthermore, we also demonstrate that our proposed spectral method is scalable and easily tunable, besides being easy to implement. Finally, we compare our method to other traditional spectral techniques for segmentation and observe that it outperforms them in terms of runtime and quality.
The next sections are organized as follows: In Section 2, we show an overview of Normalized Cuts, its main algorithm, and drawbacks. In Section 3, we present our proposed graph construction and characterize the solution of Normalized Cuts when optimized to it. Experimental results are presented in Section 4 along with a brief discussion on them. We finish with conclusions and future work in Section 5.
2 Background
2.1 Normalized Cuts
Let $I$ be an image of $n$ pixels and $k$ unique colors. Let $G=(V,E,w)$ be an undirected weighted graph where each node corresponds to a pixel and the weight function $w(i,j)$ evaluates the similarity between pixels $i$ and $j$. A binary segmentation of $I$ is achieved by finding a partition $(A,B)$ of $V$ that minimizes the cut value $\operatorname{\operatorname{Cut}}(A,B)=\sum_{i\in A,j\in B}w(i,j)$.
This approach however favors unbalanced partitions, leading to undesirable segmentations. One successful way of addressing this issue is to consider the normalized cut value $\operatorname{\operatorname{NCut}}(A,B)$ [3] instead,
$\operatorname{\operatorname{NCut}}(A,B)=\frac{\operatorname{\operatorname{Cut}}(A,B)}{\operatorname{\operatorname{Vol}}(A)}+\frac{\operatorname{\operatorname{Cut}}(A,B)}{\operatorname{\operatorname{Vol}}(B)}.$
Here $\operatorname{\operatorname{Vol}}(A)$ is a measure of the “volume” of $A$ defined as $\operatorname{\operatorname{Vol}}(A)=\sum_{i\in A,j\in V}w(i,j)$. Later, we use $\operatorname{\operatorname{Vol}}(V)=\operatorname{\operatorname{Vol}}(A)+\operatorname{\operatorname{Vol}}(B)$ to represent the volume of all edges in $G$.
Let $W$ be the weighted adjacency matrix of $G$ and $D$ be the diagonal degree matrix with $D(i,i)=\sum_{j\in V}w(i,j)$. The Laplacian matrix of $G$ is then defined as $L=D-W$. Following the derivation presented in [15], we can write the problem of minimizing the $\operatorname{\operatorname{NCut}}$ of $G$ as:
$\min_{\mathbf{x}}\quad$ $\mathcal{E}(\mathbf{x}|G),$ (1)
s.t. ${\mathbf{x}^{\top}D\mathbf{x}}=1,\quad\mathbf{1}^{\top}D\mathbf{x}=0,\quad\mathbf{x}\in\left\{-\alpha,\beta\right\}^{n}.$
where $\mathbf{x}$ is the vector that corresponds to our final segment assignment, $\mathcal{E}(\mathbf{x}|G)=\mathbf{x}^{\top}L\mathbf{x}$ is the objective function or assignment energy we wish to minimize and:
$\alpha=\sqrt{\frac{\operatorname{\operatorname{Vol}}(A)}{\operatorname{\operatorname{Vol}}(V)\operatorname{\operatorname{Vol}}(B)}},\quad\beta=\sqrt{\frac{\operatorname{\operatorname{Vol}}(B)}{\operatorname{\operatorname{Vol}}(V)\operatorname{\operatorname{Vol}}(A)}}.$ (2)
The algorithm in [3], here called the $\operatorname{\operatorname{NCut}}$ algorithm, solves a relaxation of the minimum $\operatorname{\operatorname{NCut}}$ problem by dropping the integer constraint in Eq. 1. The new optimization problem can be solved via generalized eigenvector decomposition,
$L\mathbf{x}=\lambda D\mathbf{x}.$ (3)
The algorithm then selects the eigenvector $\mathbf{x}$ with second-smallest eigenvalue, and partitions $V$ by thresholding $\mathbf{x}$ [3].
2.2 Classical Edge Weight for Image Segmentation
A crucial step in developing efficient segmentation solvers using $\operatorname{\operatorname{NCut}}$ is to define how the function $w$ is computed. In the original $\operatorname{\operatorname{NCut}}$ work and in early subsequent work [16, 17, 18], the $w$ combined two grouping cues in a single value:
[…]
×$\times\mathds{1}(\lVert X(i)-X(j)\rVert<r),$
where $I(j)$ and $X(j)$ are the appearance (such as the color) of pixel $j$ and its spatial location on image $I$, respectively, $\mathds{1}(\cdot)$ is the indicator function, and $r$, $\sigma_{I}$ and $\sigma_{X}$ have fixed values.
2.3 Drawbacks of the Traditional Approach
The $\operatorname{\operatorname{NCut}}$ formulation based on the above weight function presents several drawbacks. For one, it is computationally inefficient to construct the above graph for high resolution images, since it requires visiting each pixel and their neighboring locations withing a radius $r$. Secondly, this graph is potentially dense, which poses memory and runtime difficulties during eigenvector decomposition. $\operatorname{\operatorname{NCut}}$ using the weight as in Eq. 4 is also known to be inefficient, even for small images [10].
Moreover, calibrating the values for $\sigma_{X}$ and $\sigma_{I}$ is challenging [14]. Finally, the optimization in Eq. 1 on the graph constructed using Eq. 4 lacks interpretability [14]. The only understanding we can extract from its solution is that it encourages similar pixels to be grouped together, which is commonplace for any spectral clustering algorithm.
3 Proposed Methodology
3.1 New Graph Formulation
Inspired by [4] and [2], we show that a sparse graph formulation efficiently tackles the issues with the classical $\operatorname{\operatorname{NCut}}$ formulation. Our graph construction starts by considering a grid graph $G_{\text{grid}}$ whose vertices correspond to image pixels, with edges of weight $\mu>0$ between each pair of neighboring pixels. From $G_{\text{grid}}$, we create a new graph $G_{\text{all}}$ by adding $k$ extra nodes, each corresponding to one of the $k$ unique colors in $I$. Each node on $G_{\text{grid}}$ is then connected to the extra node corresponding to its color with an edge of unitary weight. This graph formulation is computationally attractive for its sparsity. Assuming $k\ll n$, $G_{\text{all}}$ only requires $O(n)$ edges. When segmenting $I$, we only make use of the assignment vector for the grid nodes.
3.2 Interpretation of the Cut on $G_{\text{all}}$
Let $W_{\text{grid}}$ and $L_{\text{grid}}$ be the adjacency and Laplacian matrices of $G_{\text{grid}}$, respectively. Let $W_{\text{all}}$ and $L_{\text{all}}$ be the analogous to $G_{\text{all}}$. These matrices are related to each other as follows:
$W_{\text{all}}=\begin{bmatrix}W_{\text{grid}}&H\\ H^{\top}&0\end{bmatrix},\quad L_{\text{all}}=\begin{bmatrix}L_{\text{grid}}+I&-H\\ -H^{\top}&D_{n}\end{bmatrix}$
where $H\in\mathbb{R}^{n\times k}$ is a one-hot encoding of the colors in $I$, i.e. $H(i,j)=\mathds{1}\{I(i)=j\}$, and $D_{n}\in\mathbb{R}^{k\times k}$ is diagonal such that $D_{n}(j,j)=n_{j}$, where $n_{j}$ is the number of pixels in $I$ with color $j$. Let $\mathbf{x}$ and $\mathbf{y}$ be the assignment vectors for the nodes in $G_{\text{grid}}$ and for the extras nodes, respectively. Note that $\mathbf{x}\in\{-\alpha,\beta\}^{n}$ and $\mathbf{y}\in\{-\alpha,\beta\}^{k}$, where $\alpha$ and, $\beta$ are computed using Eq. 2 on the graph $G_{\text{all}}$. Let $\mathbf{z}=[\mathbf{x}|\mathbf{y}]$ be $\mathbf{x}$ and $\mathbf{y}$ concatenated. The objective in Eq. 1 then becomes:
$\mathcal{E}(\mathbf{z}|G_{\text{all}})$ $=\mathbf{z}^{\top}L_{\text{all}}\mathbf{z}$ (5)
$=\mathbf{x}^{\top}(L_{\text{grid}}+I)\mathbf{x}-2(H\mathbf{y})^{\top}\mathbf{x}+\mathbf{y}^{\top}D_{n}\mathbf{y}.$
Noticing that $D_{n}=H^{\top}H$, we derive the energy associated with the assignment $\mathbf{z}$ of the nodes on $G_{\text{all}}$:
$\mathcal{E}(\mathbf{z}|G_{\text{all}})=\lVert\mathbf{x}-H\mathbf{y}\rVert^{2}_{2}+\mathbf{x}^{\top}L_{\text{grid}}\mathbf{x}.$ (6)
Table 1: Quantitative results on synthetic images in terms of JAC and runtime. The data was generated using four different noise models on three different ground truth patterns. The lower the JAC value, the better.
[…]
tab_synt
Original
Proposed
NCut
MixNCut
MGSC
Fig. 1: Qualitative results on synthetic images. Top: Gaussian noise, $\sigma_{g}^{2}=1.5$. Bottom: Salt & Pepper, $d=0.7$.
The first term in the above expression accounts for the number of pixels that are assigned to a partition that is not the same as the partition its color is assigned to. In fact, if we define $n_{\text{missmatch}}$ as the number of these mismatches, we have:
$\lVert\mathbf{x}-H\mathbf{y}\rVert^{2}_{2}=(\alpha+\beta)^{2}n_{\text{missmatch}}.$ (7)
Now, because of the grid nature of $G_{\text{grid}}$, the second term in Eq. 6 accounts for the number of neighboring pixel pairs that are assigned to different segments:
$\mathbf{x}^{\top}L_{\text{grid}}\mathbf{x}=\mu(\alpha+\beta)^{2}n_{\text{boundary}},$ (8)
where $n_{\text{boundary}}$ stands for the number of pixels on the boundary of the final segmentation. Finally, noticing that $(\alpha+\beta)^{2}=\operatorname{\operatorname{Vol}}(V)/(\operatorname{\operatorname{Vol}}(A)\operatorname{\operatorname{Vol}}(B))$ from Eq. 2, we have that:
$\mathcal{E}(\mathbf{z}|G_{\text{all}})=\frac{\operatorname{\operatorname{Vol}}(V)}{\operatorname{\operatorname{Vol}}(A)\operatorname{\operatorname{Vol}}(B)}(n_{\text{missmatch}}+\mu n_{\text{boundary}}).$ (9)
Minimizing $\mathcal{E}(\mathbf{z}|G_{\text{all}})$ for $\mathbf{x}$ and $\mathbf{y}$, results in finding a partition of $G_{\text{all}}$ that trades off between spacial coherence (neighboring pixels should belong to the same segment) and color cohesion (pixels in a segment should have colors belonging to the same subset of available intensities) via the value of $\mu$. Minimizing the multiplicative factor in Eq. 9 also encourages a balanced partition of the nodes in $G_{\text{all}}$.
(a) Image
(b) $\mu=5$ ($8.63s$)
(c) $\mu=50$ ($23.22s$)
(d) $\mu=300$ ($104.56s$)
Fig. 2: Qualitative results on RBG images for various $\mu$. (b)-(d) show the eigenvector from Eq. 3 (left) and the corresponding segmentation (right). The image sizes are $800\times 618$ (top) and $800\times 524$ (bottom). Average runtimes are shown in parentheses.
3.3 Improvements
3.3.1 Soft color linkage
Despite the simplicity and computational efficiency of the above result, it does not explicitly use the pixel color intensities. This means that we would still achieve the same solution to Eq. 1 even if we randomly relabeled the set of available colors. In our experiments, we test the performance of a softer version of the graph described in the previous sections. In this new graph construction, each node $i$ in $G_{\text{grid}}$ is linked to all extra nodes with an edge of weight given by
$w^{\prime}(i,c)=\exp\left(\frac{||I(i)-c||^{2}}{2\sigma^{2}}\right)$ (10)
instead of the unitary weight from the original approach. The interpretation of the solution of Eq. 1 under this new graph is left for future work.
3.3.2 Case when $k$ is large
Our method assumes $k$, the number of unique colors in $I$, to be small compared to $n$. Although this assumption is satisfied for gray scale images, it becomes unrealistic for colored ones. In this case, we vectorize the RGB color space using $K$-means clustering. The weights in Eq. 10 are then computed using the average RGB color in each cluster. This approach also alleviates the issue posed in the previous section.
4 Numerical Experiments
4.1 Experimental Setup
For performance evaluation, we assessed the segmentations via the Jaccard Dissimilarity measure (JAC), defined as:
$\text{JAC}=1-\max(J(R,S),J(R,\neg S))$ (11)
where $R$ and $S$ are the ground truth image and segmentation, respectively, and $J(R,S)=|R\cap S|/|R\cup S|$ is the Jaccard Index [19]. $\neg S$ is the segmentation with labels swapped.
For our synthetic results, we contaminate three different background and foreground square patterns (intensities of either 0 and 1) of side length $\ell$ with either additive zero mean and variance $\sigma_{g}^{2}$ Gaussian noise or Salt & Pepper with noise density $d$. The generated images are then rescaled and discretized, so the color intensities are integers in $[0,255]$. We also test our method using two RGB images from the Alpha Matting dataset [20]. For each of them, we set $k=256$ in their preprocessing stage (ref. Section 3.3.2).
We compare our proposed method with $\operatorname{\operatorname{NCut}}$ and two more spectral methods for unsupervised segmentation: Multi Graph Spectral Clustering (MGSC) [21] and MixNCut [14]. For the later two, we create two separate graphs, a grid one for pixel locations and a fully connected one for color data, as described in [14]. When using all these techniques, we refrain from any sparsification algorithm for simplicity. For $\operatorname{\operatorname{NCut}}$, we also set $r=\infty$ in Eq. 4 for experimental simplicity and so that we can compare similar graph formulations.
The experiments ran on an Apple M1 Pro Chip (8 core CPU, 14 core GPU) with 16 Gb of RAM memory and were coded in Python 3.9.7 using Numpy 1.21.4 and Scipy 1.7.3111A demo code for this paper is available at www.github.com/rahul-palnitkar/Sparse-Graph-Spectral-Segmentation..
4.2 Results and Discussion
In Table 1, we show the quantitative assessment of our proposed method for images with $\ell=100$ pixels and two different pairs of noise contaminations for generating each synthetic image under the three proposed ground-truth patterns. For our method, we depict the best JAC when considering the result from trying $\mu$ and $\sigma\in[0,1,\ldots,10]$. For the other methods, we tuned their available parameters and report their best JAC results. In the same table, we also display the average runtime for generating each of the best results’ performance in each method per pattern. In Figure 1, we qualitatively compare the best results of each method under two challenging noise levels. These results demonstrate that our method greatly over performs the comparative methods in terms of runtime, a consequence of its sparse nature, while also being competitive in terms of JAC, attaining almost perfect segmentations in most challenging scenarios.
In Figure 2, we qualitatively evaluate the effect of $\mu$ to the final segmentation. In these experiments, $\sigma$ is set to 0. In this same figure, we also qualitatively demonstrate our method’s performance and runtime in large RGB images. Here, we note that $\mu$ has the effective role of smoothing the borders of the final segmentation, removing artifacts or boundary details as it increases.
Finally, in Figure 3 we present the average runtime and JAC values for the segmentations of the three patterns of them. We vary the image side lengths $\ell$ from $100$ to $1000$ and contaminated each image with zero mean Gaussian noise with $\sigma_{g}^{2}=\ell/100$, which means that images get increasingly hard with size. We again keep $\sigma=0$ in our proposed algorithm. These results demonstrate that our method scales well despite its spectral nature and generates high quality segmentations on large images even in very challenging noise settings.
5 Conclusion and Future Directions
In this paper, we introduce a new spectral algorithm for image segmentation that overcomes some historical issues found in the Normalized Cuts framework and in its related methods. We employ an efficient sparse graph, whose usage under the traditional Normalized Cuts optimization problem leads to an interpretable cut value. We also show that our method is fast and produces high quality results for large RGB imagery and under high synthetic noise settings. For future work, we will derive an interpretation of the cut found on the graph proposed in Section 3.3.1 and extend our methodology to superpixel images—which can further still improve the overall scalability of our proposed method—and multi-region image segmentation.
[ scale only axis, height=1.5cm, width=.75ytick pos=left, ylabel=JAC, yticklabel style = font=, color=red, xticklabel style = font=, every y tick/.style=red, ylabel near ticks, xlabel near ticks, xmin=100,xmax=800, ymax=0.1, scaled y ticks=true, yticklabel style=/pgf/number format/fixed ]
[mark=none, red] table[x index=0,y index=2] data/size_times_average_variable.dat;
[…]
Fig. 3: Segmentation results for varying image sizes. Left: a sample image for $\ell=500$. Right: the runtime and JAC value results varying $\ell$ (on the $x$-axis).